# Behavioral alignment in social networks

### v1

@@ Abstract
{adjustwidth}
1cm1cm The orderly behaviors observed in large-scale groups, such as fish schooling and the organized movement of crowds, are both ubiquitous and essential for the survival and stability of these systems. Such complex collective behaviors often emerge from simple local interactions and strategy adjustments among individuals. Understanding how these basic rules shape complex group dynamics has long been a significant scientific challenge. Historically, research has predominantly focused on imitation and social learning, where individuals adopt the strategies of more successful peers to refine their behavior. However, in recent years, an alternative learning approach—self-exploration and introspective learning—has garnered increasing attention. In this paradigm, individuals assess their own circumstances and select strategies that best align with their specific conditions. Two primary forms of this learning are coordination and anti-coordination, where individuals align with and diverge from the local majority, respectively. In this study, we analyze networked systems of coordinating and anti-coordinating individuals, exploring the combined effects of system dynamics, network structure, and behavioral patterns. We address several practical questions, including the number of equilibria, their characteristics, the equilibrium time, and the resilience of systems. We find that the number of equilibrium states can be extremely large, even increasing exponentially with minor alternations to the network structure. Moreover, the network structure has a significant impact on the average equilibrium time. Despite the complexity of these findings, variations can be captured by a single, simple network characteristic: the average path length. Our research offers valuable insights into how modifications to the interaction structure can influence behavioral alignment in social networks.
@@ Introduction
The nature of individual decision-making processes can profoundly influence evolutionary outcomes in large-scale systems, including many populations found in social, biological, economic, and ecological contexts [1, 2, 3, 4]. Imitation is an important and well-studied social learning mechanism, in which individuals copy the behavior of those who are more successful or authoritative [5, 6]. When combined with social network structure, imitation can drive the evolution of prosocial traits like cooperation [7], behaviors which otherwise would not be chosen by perfectly rational agents. A drawback of imitation is that it requires individuals to have extensive information, including the behaviors of their direct neighbors as well as the neighbors of those neighbors, both to assess the behaviors and imitate those they evaluate as being successful [8].
@@
Conversely, coordination and anti-coordination as decision-making approaches allow individuals to independently seek the most beneficial strategy based on analyzing information available from interaction partners. Coordination, where individuals conform to the majority, and anti-coordination, where individuals diverge from the majority, can be found in many real-life scenarios. Examples of coordination include the formation of driving conventions, where people spontaneously drive on the same side of the road, and the choice of language within multilingual communities, where individuals tend to use the language adopted by most people [9, 10]. On the other hand, instances of anti-coordination include auction strategies, where some bidders choose less popular items in order to avoid competition, and problems of resource allocation, where variation in resource consumption and use can mitigate over-exploitation [11, 12].
@@
It is known that systems following either coordinating or anti-coordinating patterns must converge to an equilibrium state under asynchronous activation [13]. In fact, convergence to an equilibrium in such systems occurs in a finite number of strategy updates [14]. However, convergence properties of the system are sensitive to initial strategy composition and behavioral switching threshold [15, 16, 17, 18]. These studies focus mainly on convergence itself, leaving some critical aspects yet to be explored. Dynamical systems often exhibit a multitude of equilibrium states, a phenomenon observed in epidemiology, network synchronization, and the evolution of ecosystems [19, 20, 21, 22]. The variety of equilibrium states can have profound implications for a system. For example, the shallow lake system has two equilibrium states: clear and turbid. The clear state is rich in submerged vegetation, while in the turbid state, animal diversity is lost and algae is reduced [23]. It is natural, then, to be concerned with a systems’ equilibrium probability distribution and equilibration time. For instance, in evolutionary game theory, fixation probabilities and times are important indicators of whether cooperation is favored [24, 25, 26, 27, 28]. We aim to fully explore these properties of coordination and anti-coordination systems in the paper.
@@
The physical structure of a system profoundly influences the outcomes of evolution [29, 30, 31, 32, 33, 34]. Degree distribution heterogeneity and the clustering coefficient represent two important measures of heterogeneity in networked systems [35, 36]. In many networks, there exist highly-connected nodes, called hubs, and nodes tend to form tightly-knit groups in networks with high clustering coefficients. Both of these properties greatly influence the dynamical characteristics of networked systems. For instance, in the disease transmission model, heterogeneous networks might see rapid outbreaks of epidemic because hubs can accelerate disease spread, and networks with high clustering coefficients may lead to intense local outbreaks, as infections spread rapidly within closely interconnected groups [37, 38]. In quite a different setting, heterogeneous networks are strongly resilient to random network attacks but remain vulnerable to targeted strategies, and networks with high clustering coefficients possess a higher percolation threshold, which increases the cost of attacks [39, 40]. Hence, thoroughly analyzing the impact of network structure is essential for advancing our understanding of network dynamics.
We begin by analyzing the number of equilibrium states, which we discover can be unexpectedly large both in coordination and anti-coordination systems. Our findings show that the degree distribution and clustering coefficient greatly influence the number of equilibrium states. In fact, the number of equilibrium states can grow by orders of magnitude with only minor changes in the network structure. Furthermore, the number of equilibrium states and the equilibration time can both be characterized by a simple measure: the average path length. As the average path length increases, the average equilibrium time increases in coordination systems and decreases in anti-coordination systems. Finally, we explore differences among these equilibrium states.
@@
We consider a population of $N$ individuals, denoted by $\mathcal{N}=\left\{1,2,\dots,N\right\}$. The spatial structure of the population is represented by a network with $N$ nodes, in which each individual occupies exactly one node, and edges represent connections between individuals. For nodes $i$ and $j$, we denote by $k_{ij}$ the edge weight between $i$ and $j$, which is $1$ if there is an edge between $i$ and $j$ and is 00 otherwise. Every individual has one of two strategies, $A$ or $B$, and plays a game with each neighbor. When both players adopt strategy $A$, they both get a payoff $a$. Similarly, when both players employ strategy $B$, they both receive a payoff $d$. If one player chooses strategy $A$ and the other uses strategy $B$, the $A$-individual gets $b$ and the $B$-individual gets $c$.
@@
In a coordination game, $a-b-c+d>0$. Thus, under best-response dynamics, Eq. 2 gives
$s_{i}\left(t+1\right)$ =$=\begin{cases}1&n_{i}>\tau k_{i},\\ s_{i}\left(t\right)&n_{i}=\tau k_{i},\\ 0&n_{i}<\tau k_{i},\end{cases}$ (3)
where, again, $\tau=\left(d-b\right)/\left(a-b-c+d\right)$, which coincides with the probability of playing $A$ in the mixed-strategy Nash equilibrium. On the other hand, $a-b-c+d<0$ in an anti-coordination game, which yields a next-round action of
$s_{i}\left(t+1\right)$ =$=\begin{cases}1&n_{i}<\tau k_{i},\\ s_{i}\left(t\right)&n_{i}=\tau k_{i},\\ 0&n_{i}>\tau k_{i}.\end{cases}$ (4)
@@
Figure 1: Strategy evolution under coordination and anti-coordination decision-making patterns. a, The structure is described by a network, where each individual (node) adopts either strategy $A$ (red) or $B$ (blue) to interact with neighbors. An individual (marked by “?”) is chosen to update its strategy to maximize payoff, interacting with $n$ individuals using strategy $A$ and $k-n$ individuals using strategy $B$. b, An example of a coordination game, where the chosen individual receives a higher payoff when adopting the strategy used by the majority of his neighbors. For this payoff matrix, taking the strategy $A$ used by the majority brings a higher payoff than taking $B$ whenever $n>k/2$. The individual opts for strategy $B$ when $n<k/2$. The individual maintains its current strategy if $n=k/2$, which is equivalent to $u^{A}=u^{B}$. c, An example of an anti-coordination game. When $n>k/2$, adopting strategy $B$ yields a higher payoff. If $n<k/2$, strategy $A$ leads to a higher payoff. The payoffs are the same when $n=k/2$.
It is worth noting what best-response dynamics would lead to for other classes of games, such as social dilemmas. In a prisoner’s dilemma, which is defined by $c>a>d>b$, we have $d-b>0$, but $a-b-c+d$ can be positive, negative, or zero. However, in all three cases, the payoff difference in Eq. 2 is negative, which means that a player’s best response is always $B$. This finding is not surprising, given that $B$ is a dominant action in a prisoner’s dilemma and will always be used by a rational agent. However, there are also weaker social dilemmas such as snowdrift games, which are defined by $c>a>b>d$. This, in fact, is an example of an anti-coordination game (just not a pure anti-coordination game). For a snowdrift game, we have $\tau\in\left(0,1\right)$; in particular, unlike in a prisoner’s dilemma, the best response is non-trivial (Eq. 4).
Results
In an equilibrium state, $s_{i}\left(t+1\right)=s_{i}\left(t\right)$ for all $i\in\mathcal{N}$. We first explore the number of equilibrium states on representative networks and conduct a quantitative analysis of a class of networks to provide intuition about the interplay between network structure and evolutionary dynamics. Next, we investigate the absorption capacity of the equilibrium states and the average time to reach them. Finally, we examine the robustness of various network structures.
@@ Results
In Fig. 2a-c, we show the number of equilibrium states as a function of the clustering coefficient. As the clustering coefficient increases, there is a decrease in the number of equilibrium states in coordination games, while this number increases in anti-coordination games. For coordination games in networks with high clustering coefficients, if interconnected nodes choose the same strategy, then it becomes highly probable for their common neighbors to adopt the strategy, eventually resulting in the strategy being employed by the vast majority or even all nodes. Therefore, the number of equilibrium states in networks with high clustering coefficients is typically lower than that in networks with low clustering coefficients. Consequently, the number of equilibrium states in coordination games is low because the entire network essentially forms a large cluster. In contrast, for anti-coordination games in networks with high clustering coefficients, individuals choose strategies opposite to most of their neighbors. Therefore, if the strategy proportions within the clusters remain at the same level, more individuals can choose between two strategies, culminating in a greater number of equilibrium states (see Fig. S2 in the Supplementary Information for an example of the effect of clustering coefficients under different games). Thus, clusters promote the spread of strategies in coordination games but act as a barrier in anti-coordination games. In other words, clusters are usually taken over by one strategy in coordination games, while they always include two opposing strategies in anti-coordination games.
Figure 2: A minor change in network structure can dramatically change the number of equilibrium states. We consider two important network characteristics: closed triads (measured by the clustering coefficient, $\mathcal{C}$) and node degree heterogeneity distribution (measured by the distribution heterogeneity coefficient, $\gamma$). We generate networks with varying clustering coefficients by rewiring edges in a degree-preserving manner, starting with regular ring lattice networks [45]. Networks with different heterogeneity coefficients are generated based on the connection kernel and the preferential attachment mechanism [36, 46]. The number of equilibrium states is presented for a broad range of clustering coefficients, $\mathcal{C}$, (a-c) and heterogeneity coefficients, $\gamma$, (d-f), for three thresholds, $\tau=0.26,0.37,0.49$, and two types of games, coordination (red) and anti-coordination (blue). For each kind of network, we sample 50 networks and present the median, quartiles, and outliers. Modifications to network structures significantly change the number of equilibrium states. Parameters: $N=30$ and average degree $\bar{k}=4$.
@@
Both the clustering coefficient and degree distribution are fundamental properties of complex networks, and can have strong impacts on evolutionary dynamics. An important quantity in the graphs shown in Fig. 2 is the average path length, which represents the mean number of edges that nodes in the network have to pass through to reach the most distant node. By reorganizing the data presented in Fig. 2 and modifying the horizontal axis to represent the average path length, we obtain Fig. 3. This figure shows that as the average path length increases, the number of equilibrium states increases, peaks, and then decreases in coordination games (see Fig. 3a-c). On the other hand, anti-coordination games see continual growth in the number of equilibrium states as the average path length of the network grows (see Fig. 3d-f).
@@
Figure 3: The number of equilibrium states as a function of average path length. Panels ad, be, and cf correspond to $\tau$ values of $0.26$, $0.37$, and $0.49$, respectively. We investigate networks with different structures, including those with variations in clustering coefficients (purple dots, corresponding to Fig. 2a-c) and in degree distribution heterogeneity (green dots, corresponding to Fig. 2d-f). Every dot is the number of equilibrium states of a network. For a given type of game, the average path length describes the number of equilibrium states: a moderate average path length leads to a higher number of equilibrium states in coordination games (a-c), whereas in anti-coordination games, a greater average path length results in a higher number of equilibrium states (d-f). Parameter values: network size $N=30$ and average degree $\bar{k}=4$.
@@
To better understand this observation, we can analyze a sequential root-leaf structure, a representative network on which it is possible to more explicitly characterize the number of equilibrium states. This network is parameterized by $n$, the number of hubs, and $m$, the number of leaves per (interior) hub (the terminal nodes have $m+1$ leaves (see Fig. 4a). For a fixed network size, increasing the number of hubs leads to a longer mean path length (but requires a decrease in the number of leaves per hub to maintain constant size). In both coordination and anti-coordination games, we provide an explicit formula for the number of equilibrium states, as a function of the number of root nodes, $n$, leaf nodes, $m$, and the behavioral switching threshold, $\tau$. The basic structure of the analysis is given in Supplementary Information, §2, where the number of equilibrium states is characterized in terms of the eigenvalues of a $4\times 4$ binary matrix.
@@
Figure 4: A sequential root-leaf structure. a, In this idealized structure, $n$ root nodes are sequentially connected. Each root node is linked to $m$ leaf nodes, with the exception that the terminal root nodes at both ends each connect to $m+1$ leaf nodes. b, In coordination games, there is an initial increase followed by a decrease in the number of equilibrium states as $n$ grows. Conversely, in anti-coordination games, the quantity of equilibrium states grows with $n$. c, Typical equilibrium states for the sequential root-leaf structure with different root nodes in coordination games. d, Typical equilibrium states for the sequential root-leaf structure with different root nodes in anti-coordination games. Red (blue) nodes in cd denote individuals of type $A$ ($B$). Parameters: network size $N=50$ and behavioral switching threshold $\tau=0.37$.
@@
In anti-coordination games, the number of equilibrium states is determined by hubs as well. We note an upward trend in the equilibrium state count with low and increasing values of $n$. But unlike the coordination game, more equilibrium states exist when $n$ is significantly large, corresponding to the cases when $m=0$ and $\tau\neq 1/2$ (see Fig. 4b). For example, when $n=48$, $m=0$, and $\tau=0.37$, there are $1{,}221{,}537$ equilibrium states. In fact, hubs vanish in networks with long path lengths, individuals must adopt strategies opposite most of their neighbors (Fig. 4d), thereby introducing more potential equilibria.
@@
Fig. 5 illustrates the degree distribution when the clustering coefficient is $0.05$ and $0.45$ and the heterogeneity coefficient is $-4$ and $4$. (See Fig. S5 in Supplementary Information for the distribution under other parameter values.) In the coordination game, there exist equilibrium states with strong absorption capacity in any network. The behavioral switching threshold determines if one strategy is favored over the other: when $\tau<0.5$, strategy $A$ is favored in coordination games and disfavored in anti-coordination games. From the data, we find that after half of the individuals in the system adopt the favored strategy, the system tends to evolve into a state where all individuals adopt the same strategy (see Fig. S6 in Supplementary Information).
Figure 5: Preferred decision-making patterns in coordination and anti-coordination games. We analyzed the arrival probability of each equilibrium state for both coordination games (a-d) and anti-coordination games (e-f). In each histogram of panel a-h, the horizontal axis represents the arrival probability, while the vertical axis indicates the number of equilibrium states whose probabilities fall within the corresponding interval. The equilibrium state with the strongest absorption capacity corresponding to each parameter is shown in each panel, where the red (blue) nodes represent individuals with strategy $A$ ($B$). Each histogram displays the results from simulations conducted $3\times 10^{7}$ times on each of $10$ distinct networks. Parameter values: network size $N=30$, average degree $\bar{k}=4$, and behavioral switching threshold $\tau=0.37$.
@@
In addition to which states the system can reach, an important quantity is the time it takes to reach these states. In this section, we consider the equilibrium time, which is measured by the number of strategy changes in the system prior to reaching an equilibrium state.

### v2

@@ ABSTRACT
The orderly behaviors observed in large-scale groups, such as fish schooling and the organized movement of crowds, are both ubiquitous and essential for the survival and stability of these systems. Understanding how such complex collective behaviors emerge from simple local interactions and behavioral adjustments is a significant scientific challenge. Historically, research has predominantly focused on imitation and social learning, where individuals adopt the strategies of more successful peers to refine their behavior. However, in recent years, an alternative learning approach based on self-exploration and introspective learning has garnered increasing attention. In this paradigm, individuals assess their own circumstances and select strategies that best align with their specific conditions. Two examples are coordination and anti-coordination, where individuals align with and diverge from the local majority, respectively. In this study, we analyze networked systems of coordinating and anti-coordinating individuals, exploring the combined effects of system dynamics, network structure, and behavioral patterns. We address several practical questions, including the number of equilibria, their characteristics, the equilibrium time, and the resilience of the system. We find that the number of equilibrium states can be extremely large, even increasing exponentially with minor alterations to the network structure. Moreover, the network structure has a significant impact on the average equilibrium time. Despite the complexity of these findings, we find that variations can be captured by a single, simple network characteristic (the average path length), which we illustrate in both synthetic and empirical networks.
Keywords: game theory, complex network, best-response, evolutionary dynamics
@@ INTRODUCTION
The nature of individual decision-making processes can profoundly influence evolutionary outcomes in large-scale systems, including many populations found in social, biological, economic, and ecological contexts [1, 2, 3, 4]. A long-standing question across these domains is how simple local interactions, collectively give rise to organized, system-level behavior, such as consensus, polarization, or innovation diffusion [5, 6]. Addressing this question requires understanding both the behavioral rules individuals follow and how these rules interact with the underlying network structures. Imitation is an important and well-studied social learning mechanism, in which individuals copy the behavior of those who are more successful or authoritative [7, 8]. When combined with social network structure, imitation can drive the evolution of prosocial traits like cooperation [9], behaviors which otherwise would not be chosen by perfectly rational agents. A drawback of imitation is that it requires individuals to have extensive information, including the behaviors of their direct neighbors as well as the neighbors of those neighbors, both to assess the behaviors and imitate those they evaluate as being successful [10].
@@
Conversely, coordination and anti-coordination as decision-making approaches allow individuals to independently seek the most beneficial strategy based on analyzing information from interaction partners. These forms of decision-making involve the ability to self-assess, enabling a deeper level of autonomy and robustness not achievable through imitation alone. Coordination, where individuals conform to the majority, and anti-coordination, where individuals diverge from the majority, can also be found in many real-life scenarios. Both are fundamental behavioral mechanisms consistently observed across diverse empirical systems. Examples of coordination include the formation of driving conventions, where people drive on the same side of the road, and the choice of language within multilingual communities, where individuals tend to use the language adopted by most people [11, 12]. On the other hand, instances of anti-coordination include auction strategies, where some bidders choose less popular items in order to avoid competition, and problems of resource allocation, where variation in resource consumption and use can mitigate over-exploitation [13, 14].
@@
It is known that systems following either coordinating or anti-coordinating patterns must converge to an equilibrium state under asynchronous activation [15]. In fact, convergence to an equilibrium in such systems occurs in a finite number of strategy updates [16]. However, convergence properties of the system are sensitive to initial strategy composition and behavioral switching threshold [17, 18, 19, 20]. These studies focus mainly on convergence itself, leaving some critical aspects yet to be explored. Dynamical systems often exhibit a multitude of equilibrium states, a phenomenon observed in epidemiology, network synchronization, and the evolution of ecosystems [21, 22, 23, 24]. The variety of equilibrium states can have profound implications for a system. For example, shallow lake systems have two equilibrium states: clear and turbid. The clear state is rich in submerged vegetation, while in the turbid state, animal diversity is lost and algae is reduced [25]. It is natural, then, to be concerned with a system’s equilibrium probability distribution and equilibration time. For instance, in evolutionary game theory, fixation probabilities and times are important indicators of whether cooperation is favored [26, 27, 28, 29, 30].
@@
The physical structure of a system profoundly influences the outcomes of evolution [31, 32, 33, 34, 35, 36]. Degree distribution heterogeneity and the clustering coefficient represent two important measures of heterogeneity in networked systems [37, 38]. In many networks, there exist highly-connected nodes, called hubs, and nodes tend to form tightly-knit groups in networks with high clustering coefficients. Both of these properties greatly influence the dynamical characteristics of networked systems. For example, heterogeneity in the interaction frequency between pairs of individuals can facilitate the evolution of cooperation [39]. In disease transmission models, heterogeneous networks might see rapid outbreaks of epidemics because hubs can accelerate disease spread, and networks with high clustering coefficients may lead to intense local outbreaks, as infections spread rapidly within closely interconnected groups [40, 41, 42]. In quite a different setting, heterogeneous networks are strongly resilient to random network attacks but remain vulnerable to targeted strategies, and networks with high clustering coefficients possess a higher percolation threshold, which increases the cost of attacks [43, 44]. Hence, thoroughly analyzing the impact of network structure is essential for advancing our understanding of network dynamics.
In this paper, we investigate the dynamics of coordination and anti-coordination systems, including key properties such as the number of equilibrium states, equilibration time, probability distributions, and system resilience. Our analysis reveals that structural features of the underlying network, such as the degree distribution and clustering coefficient, strongly affect the number of equilibrium states. Notably, even small changes in network topology can lead to orders-of-magnitude variations in the number of equilibria. We further show that both the number of equilibrium states and the time required to reach equilibrium are governed by a simple topological measure: the average path length. As the average path length increases, coordination systems tend to equilibrate more slowly, while anti-coordination systems tend to equilibrate more rapidly. Furthermore, we validate these theoretical findings using twenty empirical networks spanning diverse domains, sizes, and densities. These include the karate club friendship network in a university setting [45], interaction patterns among Mexican political elites [46], a Facebook friendship network within a computer company [47], and social associations among wild Grévy’s zebras [48].
RESULTS
@@
We consider a population of $N$ individuals, denoted by $\mathcal{N}=\left\{1,2,\dots,N\right\}$. The spatial structure of the population is represented by a network with $N$ nodes, in which each individual occupies exactly one node, and edges represent connections between individuals. For nodes $i$ and $j$, we denote by $k_{ij}$ the edge weight between $i$ and $j$, which is $1$ if there is an edge between $i$ and $j$ and is 0 otherwise. Every individual has one of two strategies, $A$ or $B$, and plays a game with each neighbor. When both players adopt strategy $A$, they both get a payoff $a$. Similarly, when both players employ strategy $B$, they both receive a payoff $d$. If one player chooses strategy $A$ and the other uses strategy $B$, the $A$-individual gets $b$ and the $B$-individual gets $c$.
@@
Figure 1: Strategy evolution under coordination and anti-coordination decision-making patterns. (a) The structure is described by a network, where each individual (node) adopts either strategy $A$ or $B$ to interact with neighbors. An individual (marked by “?”) is chosen to update its strategy to maximize payoff, interacting with $n$ individuals using strategy $A$ and knk-n individuals using strategy $B$. (b) An example of a coordination game, where the chosen individual receives a higher payoff when adopting the strategy used by the majority of his neighbors. For this payoff matrix, taking the strategy AA used by the majority brings a higher payoff than taking $B$ whenever $n>k/2$. The individual opts for strategy $B$ when $n<k/2$. The individual maintains its current strategy if $n=k/2$, which is equivalent to $u^{A}=u^{B}$. (c) An example of an anti-coordination game. When $n>k/2$, adopting strategy $B$ yields a higher payoff. If $n<k/2$, strategy AA leads to a higher payoff. The payoffs are the same when $n=k/2$.
@@
In a coordination game, abc+d>0a-b-c+d>0. Thus, under best-response dynamics, Eq. 2 gives
$s_{i}\left(t+1\right)$ $=\begin{cases}1&n_{i}>\tau k_{i},\\ s_{i}\left(t\right)&n_{i}=\tau k_{i},\\ 0&n_{i}<\tau k_{i},\end{cases}$ (3)
where, again, $\tau=\left(d-b\right)/\left(a-b-c+d\right)$, which coincides with the probability of playing $A$ in the mixed-strategy Nash equilibrium. On the other hand, abc+d<0a-b-c+d<0 in an anti-coordination game, which yields a next-round action of
$s_{i}\left(t+1\right)$ $=\begin{cases}1&n_{i}<\tau k_{i},\\ s_{i}\left(t\right)&n_{i}=\tau k_{i},\\ 0&n_{i}>\tau k_{i}.\end{cases}$ (4)
@@
It is worth noting what best-response dynamics would lead to for other classes of games, such as social dilemmas. In a prisoner’s dilemma, which is defined by $c>a>d>b$, we have db>0d-b>0, but abc+da-b-c+d can be positive, negative, or zero. However, in all three cases, the payoff difference in Eq. 2 is negative, which means that a player’s best response is always $B$. This finding is not surprising, given that $B$ is a dominant action in a prisoner’s dilemma and will always be used by a rational agent. However, there are also weaker social dilemmas such as snowdrift games, which are defined by $c>a>b>d$. This, in fact, is an example of an anti-coordination game (just not a pure anti-coordination game). For a snowdrift game, we have $\tau\in\left(0,1\right)$; in particular, unlike in a prisoner’s dilemma, the best response is non-trivial (Eq. 9).
A configuration of $A$ and $B$ is defined as an equilibrium state if, under the update rule in Eq. 8 or Eq. 9, each individual’s prescribed next action matches their current action. A state $\mathbf{s}^{\ast}=\left(s_{1}^{\ast},\dots,s_{N}^{\ast}\right)$ is an equilibrium if $u_{i}\left(\mathbf{s}^{\ast}\right)=\max\left\{u_{i}^{A}\left(\mathbf{s}^{\ast}\right),u_{i}^{B}\left(\mathbf{s}^{\ast}\right)\right\}$ for all $i\in\mathcal{N}$, where $u_{i}\left(\mathbf{s}^{\ast}\right)$ denotes player $i$’s payoff in state $\mathbf{s}^{\ast}$, and $u_{i}^{X}\left(\mathbf{s}^{\ast}\right)$ ($X\in\left\{A,B\right\}$) is player $i$’s payoff when the other players use $s_{j}^{\ast}$ $\left(j\neq i\right)$ and $i$ adopts strategy $X$. Consequently, no individual has the incentive to change strategy in an equilibrium state and $s_{i}^{\ast}\left(t+1\right)=s_{i}^{\ast}\left(t\right)$ for all $i\in\mathcal{N}$. We first explore the number of equilibrium states on representative networks and conduct a quantitative analysis of a class of networks to provide intuition about the interplay between network structure and evolutionary dynamics. Next, we investigate the absorption capacity of the equilibrium states and the average time required to reach them. Here, we define the absorption capacity of a given equilibrium state as the probability that the system, starting from a random initial configuration, eventually reaches that state. Finally, we examine the robustness of various network structures.
@@
Figure 2: A minor change in network structure can dramatically change the number of equilibrium states. We consider two important network characteristics: closed triads (measured by the clustering coefficient, $\mathcal{C}$) and node degree heterogeneity distribution (measured by the distribution heterogeneity coefficient, $\gamma$). We generate networks with varying clustering coefficients by rewiring edges in a degree-preserving manner, starting with regular ring lattice networks [50]. Networks with different heterogeneity coefficients are generated based on the connection kernel and the preferential attachment mechanism [38, 51]. The number of equilibrium states is presented for a broad range of clustering coefficients, $\mathcal{C}$, (a-c) and heterogeneity coefficients, $\gamma$, (d-f), for three thresholds, $\tau=0.26,0.37,0.49$, and two types of games, coordination and anti-coordination. For each kind of network, we sample 50 networks and present the median, quartiles, and outliers. Modifications to network structures significantly change the number of equilibrium states. Parameters: $N=30$ and average degree $\bar{k}=4$.
@@
Here, we perform an exhaustive search of the entire state space to identify and count all possible equilibrium states. In Fig. 2a-c, we show the number of equilibrium states as a function of the clustering coefficient. As the clustering coefficient increases, there is a decrease in the number of equilibrium states in coordination games, while this number increases in anti-coordination games. For coordination games in networks with high clustering coefficients, if interconnected nodes choose the same strategy, then it becomes highly probable for their common neighbors to adopt the strategy, eventually resulting in the strategy being employed by the vast majority or even all nodes. Therefore, the number of equilibrium states in networks with high clustering coefficients is typically lower than that in networks with low clustering coefficients. Consequently, the number of equilibrium states in coordination games is low because the entire network essentially forms a large cluster. In contrast, for anti-coordination games in networks with high clustering coefficients, individuals choose strategies opposite to most of their neighbors. Therefore, if the strategy proportions within the clusters remain at the same level, more individuals can choose between two strategies, culminating in a greater number of equilibrium states (see Fig. S2 in the Supplementary Information for an example of the effect of clustering coefficients under different games). Thus, clusters promote the spread of strategies in coordination games but act as a barrier in anti-coordination games. In other words, clusters are usually taken over by one strategy in coordination games, while they always include two opposing strategies in anti-coordination games.
@@
Figure 3: The number of equilibrium states as a function of average path length. Panels ad, be, and cf correspond to $\tau$ values of $0.26$, $0.37$, and $0.49$, respectively. We investigate networks with different structures, including those with variations in clustering coefficients (corresponding to Fig. 2a-c) and in degree distribution heterogeneity (corresponding to Fig. 2d-f). Every dot is the number of equilibrium states of a network. For a given type of game, the average path length describes the number of equilibrium states: a moderate average path length leads to a higher number of equilibrium states in coordination games (a-c), whereas in anti-coordination games, a greater average path length results in a higher number of equilibrium states (d-f). Parameter values: network size $N=30$ and average degree $\bar{k}=4$.
Both the clustering coefficient and degree distribution are fundamental properties of complex networks, and can have strong impacts on evolutionary dynamics. An important quantity in the graphs shown in Fig. 2 is the average path length, which represents the mean number of edges that nodes in the network have to pass through to reach the most distant node. By reorganizing the data presented in Fig. 2 and modifying the horizontal axis to represent the average path length, we obtain Fig. 3. Here, the clustering coefficient is positively correlated with the average path length, and the heterogeneity coefficient is negatively correlated with it (see details in Fig. S24). Figure 3 shows that as the average path length increases, the number of equilibrium states increases, peaks, and then decreases in coordination games (see Fig. 3a-c). On the other hand, anti-coordination games see continual growth in the number of equilibrium states as the average path length of the network grows (see Fig. 3d-f).
@@
To better understand this observation, we can analyze a sequential root-leaf structure, a representative network on which it is possible to more explicitly characterize the number of equilibrium states. This network is parameterized by $n$, the number of hubs, and $m$, the number of leaves per (interior) hub (the terminal nodes have $m+1$ leaves (see Fig. 4a). For a fixed network size, increasing the number of hubs leads to a longer average path length (but requires a decrease in the number of leaves per hub to maintain constant size). In both coordination and anti-coordination games, we provide an explicit formula for the number of equilibrium states, as a function of the number of root nodes, $n$, leaf nodes, $m$, and the behavioral switching threshold, $\tau$. The basic structure of the analysis is given in Supplementary Information, §2, where the number of equilibrium states is characterized in terms of the eigenvalues of a $4\times 4$ binary matrix.
@@
Figure 4: A sequential root-leaf structure. (a) In this idealized structure, $n$ root nodes are sequentially connected. Each root node is linked to $m$ leaf nodes, with the exception that the terminal root nodes at both ends each connect to $m+1$ leaf nodes. (b) In coordination games, there is an initial increase followed by a decrease in the number of equilibrium states as $n$ grows. Conversely, in anti-coordination games, the quantity of equilibrium states grows with $n$. (c) Typical equilibrium states for the sequential root-leaf structure with different root nodes in coordination games. (d) Typical equilibrium states for the sequential root-leaf structure with different root nodes in anti-coordination games. Red (blue) nodes in cd denote individuals of type AA ($B$). Parameters: network size $N=50$ and behavioral switching threshold $\tau=0.37$.
@@
In anti-coordination games, the number of equilibrium states is determined by hubs as well. We note an upward trend in the equilibrium state count with low and increasing values of $n$. But unlike the coordination game, more equilibrium states exist when $n$ is significantly large, corresponding to the cases when $m=0$ and $\tau\neq 1/2$ (see Fig. 4b). For example, when $n=48$, $m=0$, and $\tau=0.37$, there are 1,221,5371{,}221{,}537 equilibrium states. In fact, hubs vanish in networks with long path lengths, individuals must adopt strategies opposite most of their neighbors (Fig. 4d), thereby introducing more potential equilibria.
@@
Figure 5: Preferred decision-making patterns in coordination and anti-coordination games. We analyzed the arrival probability of each equilibrium state for both coordination games (a-d) and anti-coordination games (e-f). In each histogram of panel (a-h), the horizontal axis represents the arrival probability, while the vertical axis indicates the number of equilibrium states whose probabilities fall within the corresponding interval. The equilibrium state with the strongest absorption capacity corresponding to each parameter is shown in each panel, where the red (blue) nodes represent individuals with strategy AA ($B$). Each histogram displays the results from simulations conducted $3\times 10^{7}$ times on each of $10$ distinct networks. Parameter values: network size $N=30$, average degree $\bar{k}=4$, and behavioral switching threshold $\tau=0.37$.
@@
Figure 5 illustrates the degree distribution when the clustering coefficient is $0.05$ and $0.45$ and the heterogeneity coefficient is 4-4 and $4$. (See Fig. S5 in Supplementary Information for the distribution under other parameter values.) In the coordination game, there exist equilibrium states with strong absorption capacity in any network. The behavioral switching threshold determines if one strategy is favored over the other: when $\tau<0.5$, strategy $A$ is favored in coordination games and disfavored in anti-coordination games. From the data, we find that after half of the individuals in the system adopt the favored strategy, the system tends to evolve into a state where all individuals adopt the same strategy (see Fig. S6 in Supplementary Information).
@@
Figure 6: Equilibrium time as a function of clustering coefficient, heterogeneity, and average path length. The average equilibrium time of different networks is shown for coordination games and for anti-coordination games. Each box in panels (a-f) represents the average equilibrium time for $50$ networks, showing the median, quartiles, and outliers. Open and solid dots in (g-i) correspond to the data in (a-c) and (d-f), respectively. Panel (adg), (beh), and (cfi) correspond to $\tau$ values of $0.26$, $0.37$, and $0.49$, respectively. The average equilibrium time increases (decreases) in coordination (anti-coordination) games as the clustering coefficient increases, while it decreases (increases) in coordination (anti-coordination) games as the heterogeneity coefficient increases. (g-i), The equilibrium time increases in coordination games and decreases in anti-coordination games as the average path length increases. The simulation is repeated $3\times 10^{7}$ times for each network. Parameter values: network size $N=30$ and average degree $\bar{k}=4$.
In addition to which states the system can reach, an important quantity is the time it takes to reach these states. In this section, we consider the equilibrium time, which is measured by the total number of strategy changes in the system prior to reaching an equilibrium state.